10 h.p moter pump out water from a well of depth 20m and files water tank of volume 2238 ltr at a hight of 10m from the ground.the running time of moter to fill the empty tank?where g=10ms-2

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Energy required to lift given amount of water to given height=mgh=\rho Vgh=Pt

P=10HP=7355W

Hencet=\dfrac{V\rho gh}{P}=\dfrac{2238L\times 1kg/L\times 10m/s^2\times 20m}{7335W}=63.7s